# A Reinforcement Learning based Motion Planner for Quadrotor Autonomous Flight in Dense Environment

### v2

@@ Abstract
Quadrotor motion planning is critical for autonomous flight in complex environments, such as rescue operations. Traditional methods often employ trajectory generation optimization and passive time allocation strategies, which can limit the exploitation of the quadrotor’s dynamic capabilities and introduce delays and inaccuracies. To address these challenges, we propose a novel motion planning framework that integrates visibility path searching and reinforcement learning (RL) motion generation. Our method constructs collision-free paths using heuristic search and visibility graphs, which are then refined by an RL policy to generate low-level motion commands. We validate our approach in simulated indoor environments, demonstrating better performance than traditional methods in terms of time span.
@@ I Introduction
Quadrotors are extensively used in a variety of applications, including rescue operations, fire and electricity inspection, and package delivery. Achieving autonomous flight in complex environments necessitates effective motion planning, which is crucial for ensuring safe and efficient navigation. State-of-the-art motion planning algorithms for quadrotors primarily focus on generating collision-free and dynamic feasible trajectories, subsequently optimizing them to achieve smoothness and aggressiveness. Time allocation plays a pivotal role in the trajectory generation process, as suboptimal time allocation can lead to inefficient trajectories and underutilization of the quadrotor’s dynamic capabilities.
However, traditional methods predominantly employ passive time allocation strategies, which may not fully exploit the quadrotor’s dynamic capabilities within its constraints. Additionally, the trajectories generated by these optimizers must be tracked by a controller, introducing delays, increased computational overhead, and discrepancies between the optimized and actual flight trajectories.
@@
To address these challenges, we propose a novel reinforcement learning (RL)-based motion planner for quadrotor autonomous flight in dense environments. Our approach aims to maximize the utilization of the quadrotor’s dynamic capabilities while relying only on a low-level controller. Our method employs a path searching algorithm based on heuristic search and visibility graphs, which constructs a collision-free, minimum-length path in a discrete grid space. This initial path is then converted into waypoints, serving as control points for the RL motion generator. The policy network generates low-level control commands based on observations of the quadrotor’s state, control points, and environmental information, optimizing the quadrotor’s motion (Fig. 1). These low-level commands are directly sent to the quadrotor’s controller, ensuring motion within predefined limits of velocity, acceleration, and attitude.
@@
Compared to existing motion planners, our method deviates from conventional trajectory generation and optimization frameworks by directly generating motion using an RL policy. It can handle discrete paths that may not satisfy kinodynamic and non-holonomic constraints, producing aggressive, collision-free motion within the specified dynamic constraints. We demonstrate the effectiveness and robustness of our method through extensive simulations and discuss its potential transferability to real-world quadrotor platforms, that will do later. Our key contributions are summarized as follows:
This letter propose a robust and efficient RL-based method that integrates visibility path searching and RL motion generation, directly generating feasible and aggressive motions from the policy.
We incorporate heuristic search and visibility graph concepts to develop a grid map-based pathfinding algorithm that generates collision-free, minimum-length paths.
We investigate optimal RL environment settings and obstacle information configurations for planning in dense environments, enabling motion generation from any initial state.
We provide extensive simulation evaluations of the proposed method and discuss its potential transferability to real-world quadrotor platforms.
@@ II Related Work
II-A Optimal Control Methods
Quadrotor motion planning has been extensively studied, with optimal control-based trajectory optimization methods prevailing as the predominant approach. These methods focus on generating smooth, dynamically feasible, and collision-free trajectories by solving optimization tasks modeled as optimal control problems. The pioneering approach is the minimum-snap trajectory generation[1] in differential flatness outputs, representing the trajectory as a polynomial and using gradient descent to iteratively modify the time segments. Several works propose two-stage solving methods where trajectory generation occurs in the second stage [2, 3, 4, 5, 6, 7]. For instance, [2] introduces a safe flight corridor composed of convex polyhedra, providing constraints for optimization. The method in [3] combines sampling-based and gradient optimization methods to achieve smooth, feasible, and safe trajectories. [4] utilizes B-spline for kinodynamic path searching, adopting an elastic optimization method to refine trajectories. The Euclidean Signed Distance Field (ESDF) map is employed in [8] to obtain collision potentials and in [5] to find a path in the velocity field, improving time allocation. [6] proposes a comprehensive pipeline for quadrotor motion planning, including kinodynamic path searching, B-spline trajectory generation, time allocation, and nonuniform B-spline optimization based on ESDF. Subsequent works [9, 10] introduce topological path and risk-aware yaw planning. Given the computational intensity of constructing ESDF, [11] proposes an ESDF-free gradient-based planner. However, [12] demonstrates a method to generate ESDF rapidly. These methods effectively generate smooth, safe, and feasible trajectories. However, they may be conservative regarding the quadrotor’s dynamic capabilities due to passive time allocation. In contrast, [13] and [14] focuses on spatial-temporal deformation, which aligns with our approach by considering both geometric and temporal planning concurrently.
@@ II-B RL based Methods
RL-based methods have gained significant traction in quadrotor motion planning due to their capability to optimize non-convex and discontinuous objectives, which traditional approaches often struggle to handle effectively. In [15], the authors propose an imitation learning (IL)-based teacher-student pipeline that directly maps noisy sensory depth images to polynomial trajectories. While IL is quite efficient, its scalability is limited by the need for expert data collection. [16] employs a topological path searching algorithm to generate guiding paths, followed by RL to generate control commands. This two-stage method performs well in selected scenarios, outperforming some traditional methods. [17] extracts features from depth images and uses a teacher-student IL framework similar to [15]. Furthermore, it generates control commands directly and includes a perception reward to guide yaw planning. In [18, 19], Song investigates the strengths and limitations of optimal control and RL, proposing that RL is advantageous as it can optimize better objectives that may be infeasible for optimal control methods to address.
Current RL-based methods often incorporate high-order dynamic parameters like body rate and attitude into their observations to compose trajectories or low-level control commands. Despite their advanced techniques, these methods frequently rely on imitation learning (IL) with expert knowledge or are tested in fixed environments, limiting their generalization and transferability to unknown environments, and failing to enable autonomous flight in new scenarios. Our RL approach also uses high-order dynamic parameters as observations but focuses on simplifying the learning tasks for RL. Instead of expecting RL to handle the entire path searching process, we use RL as an optimizer to refine motion and control points, similar to B-spline optimization schemes. This strategy allows for more efficient learning and better adaptation to complex environments.
@@ III System Overview
III-A Quadrotor dynamics
We model the quadrotor as a linear system with translational and rotational dynamics decoupled. The dynamics equations are
where is position in earth frame, is velocity, is rotation matrix from body frame to earth frame, is z axis in body frame, is thrust of each rotor, is mass, is drag coefficient, is diagonal inertia matrix, is angular velocity in body frame, and is gyroscopic torque and torque by thrusts respectively. We use Euler angles to represent rotation and set a limit on roll and pitch angles to avoid singularity.
@@ III-A Quadrotor dynamics
where denotes rotor’s torque coefficient, and represents the distance from the center of mass to the rotor center. The complete state of the quadrotor is defined as , with indicating roll, pitch and yaw angle.The control input is given by . We use simpson’s rule to do numerical integration for state update.
III-B Visibility RL Planning Structure
The proposed planning framework is illustrated in Fig. 2. The planning process operates in two stages. First, the Visibility Path Searching algorithm constructs a reference path from the start to the goal, ensuring obstacle avoidance and path length minimization. This reference path is discretized into waypoints based on the current position and the distance to the next obstacle corner (Section IV). Second, the RL Motion Generation method produces control commands to optimize the quadrotor’s motion. These commands enable the quadrotor to follow the waypoints within a certain tolerance, determining the required proximity to waypoints and strategies for obstacle avoidance in continuous space, as presented in Section V.
@@ IV Visibility Path Searching
Our pathfinding algorithm integrates the A*, Jump Point Search (JPS)[20], and Visibility Graph methods. We inherit the heuristic function and the open and closed list architecture from the A* algorithm but use a different cost function. Elements of the JPS method are incorporated to reduce the number of nodes in the open list, minimizing computational costs associated with sorting these nodes. Additionally, we leverage the optimal distance characteristics of the Visibility Graph method without constructing complex polygon edges and vertices.
The proposed strategy facilitates omnidirectional movement through obstacle corner detection and visibility checks, rather than restricting movement to increments of in 2D plane. This approach is implemented by selecting obstacle corners in the Corner function and assessing voxel visibility in the VisibleCheck function. Consequently, this method reduces the computational and memory overhead typically required for accessing each vertex and edge of the voxels. The detailed algorithm is presented in Algorithm 1.
@@ V-A Policy Framework
We construct the control points optimizing and motion generating problem in a reinforcement learning (RL) framework using Markov Decision Process (MDP) defined by tuple . The quadrotor agent begins in a state that is selected from state space distribution . In this setting, the agent interacts with the environment by sampled action from a random policy . After that, the state of the agent will be updated from to by the transition probability , and a reward will be given to the agent. Detailed configuration of the state space, action space, and reward function will be discussed in the following sections.
@@ V-A1 Observation and action space
The observation space contains three main parts: state of the quadrotor, waypoints information and ESDF based obstacle details. Specifically, we denote the state vector as
And the observation vector is derived from the state vector, incorporating some elements directly from it, including . We convert other state elements to low level features, which would be more generalized and easier for policy training. For example, waypoints position combined with absolute position of the quadrotor are used to calculate relative position of the quadrotor to target waypoints. We also add a normalized vector representing the direction from current position to the next waypoint in body frame, which is used to guide the limit of field of view (FOV) of depth camera with direction of in body frame. The detailed obstacle information is composed by the ESDF map of the environment, which will be discussed in the following section.
The action produced by RL policy is , is the collective thrust by all motors and is the desired body rate, which has been proved to perform well in learning-based quadrotor control and planning tasks. Although this control command is usually used in drone racing tasks due to its low latency and high frequency, it also takes a significant gap between simulation and real world. Despite this drawback, this approach can result in better smoothness in acceleration and velocity, which will be beneficial in energy saving and safety.
@@ V-A2 Reward function
According to paper RL vs OC, RL outperforms Optimal Control (OC) in terms of optimizing a better objective that OC may not be able to solve. From this perspective, we design a reward function that explicitly represents better active time allocation, collision avoidance, and smoothness in motion based on RL vs OC. The reward function is defined as
where is the customized progress reward, is punishment for state violation, is the reward for arriving at one waypoint, is the punishment for collision, and is the penalty for large angular velocity.
@@
We normalize the progress reward within interval and subsequently introduce a negative term , thereby imposing a time penalty which facilities time allocation and maximized the utilization of the quadrotor’s dynamic capabilities.
For state violation penalty , we impose limits on acceleration and velocity to ensure dynamic smoothness and safety. Additionally, attitude limits are set to prevent overly aggressive maneuvers and avoid singularities.
Specifically, represents the number of elements from the set that exceed their prescribed limits.
@@ V-B Training methodology
We create a training environment based on OpenAI gym, providing an interface for creating a customized environment. At the beginning of each episode, the velocity, body rate, orientation, waypoints and obstacles are randomly initialized following a predefined probabilistic pattern which is shown in Fig.3, while the position if fixed at the center of the map.
V-B1 Quadrotor state initialization
Since we are using only two waypoints for the whole navigation, we make the first waypoint totally randomized with only the distance from quadrotor position to the waypoint withing a certain range.
where is a randomly generated unit vector, and and define the waypoint distance range. The second waypoint is given by
The second waypoint is set to within a certain distance and angle from the first waypoint and the line from quadrotor position to the first waypoint.
where is a unit vector such that , with being the angle between and . must lie within the waypoint distance range and ensure that the distance between the quadrotor and the second waypoint exceeds .
Concerning velocity, constraints are applied to ensure the velocity direction does not deviate significantly, expressed as , as the path searching algorithm is expected not to produce waypoints that are in a direction completely opposite to the current velocity. We also set constraints on the orientation of the quadrotor as
V-B2 Obstacles setting
In our framework, RL is expected to generate motions based on collision-free waypoints obtained after path searching. These waypoints might still be close to or intersect with obstacles after sampling. Therefore, we only need to place obstacles on or near the lines from the quadrotor to the waypoints, rather than creating a maze-like environment. Additionally, it is essential to ensure a feasible path exists between the quadrotor position and the waypoints (see Fig. 4).
V-B3 Obstacle avoidance
Although the guided waypoints from path searching can avoid obstacles in voxel level, the quadrotor still needs to avoid obstacles in continuous space, which is achieved adding related observation and reward in RL framework.
We truncate ESDF map into TSDF map due to two reasons: first, we expect RL framework as a local motion generator to be more efficient in terms of active time allocation and collision avoidance; second, the ideal range of depth camera is limited, which is more suitable for TSDF map.
While SDF maps can address some motion planning problems, they often require the entire map or significant portions. In an RL framework, excessive information can hinder generalization. Using SDF data can aid in obstacle avoidance but is not comprehensive for all scenarios, as illustrated in Fig. 5. A more efficient method for quadrotor environmental understanding is necessary.
To address these issues and better align with our objectives, we propose a method named ray guiding, which provides understanding the basic distribution of obstacles by ray sampling and accurate velocity limitations by ray searching. To comprehensively assess the surrounding obstacles, we sample ray castings from the quadrotor’s position to the first waypoint at irregular angles centered around the waypoint direction (see Fig. 6). The end positions of each beam are used in the observation space. While this method offers a broad understanding of the space around the quadrotor, its accuracy is affected by the interval of rays, the sensing range, and the map’s resolution.
Since ray sampling provides a rough obstacle distribution, we further refine the observation using ray search. The core idea is to define safe regions in the velocity field that guide the quadrotor’s velocity towards the waypoint safely, enhancing both efficiency and reliability (see in Fig. 7). We first cast rays from the quadrotor towards the waypoints, spreading them in a cone shape to detect obstacles’ edges and free space within a certain range. If the direct ray to the waypoint is blocked, we search for a transition point from obstacle to free space. If unblocked, we identify the obstacle edges. The final ray search observation for a waypoint is given as .
V-C Sim to real transfer

### v4

# RESC: A Reinforcement Learning Based Search-to-Control Framework for Quadrotor Local Planning in Dense Environments

@@ Abstract
Agile flight in complex environments poses significant challenges to current motion planning methods, as they often fail to fully leverage the quadrotor dynamic potential, leading to performance failures and reduced efficiency during aggressive maneuvers. Existing approaches frequently decouple trajectory optimization from control generation and neglect the dynamics, further limiting their ability to generate aggressive and feasible motions. To address these challenges, we introduce an enhanced Search-to-Control planning framework that integrates visibility path searching with reinforcement learning (RL) control generation, directly accounting for dynamics and bridging the gap between planning and control. Our method first extracts control points from collision-free paths using a proposed heuristic search, which are then refined by an RL policy to generate low-level control commands for the quadrotor controller, utilizing reduced-dimensional obstacle observations for efficient inference with lightweight neural networks. We validate the framework through simulations and real-world experiments, demonstrating improved time efficiency and dynamic maneuverability compared to existing methods, while confirming its robustness and applicability.
@@ I Introduction
Quadrotors are extensively used in a variety of applications, including rescue operations, fire and electricity inspection, and package delivery [1]. Achieving autonomous flight in complex environments necessitates effective motion planning, which is crucial for ensuring safe and efficient navigation. Current state-of-the-art motion planning algorithms for quadrotors primarily focus on generating collision-free and dynamic feasible trajectories, subsequently optimizing them to achieve smoothness and aggressiveness [2]. Time allocation plays a pivotal role in the trajectory generation process, as suboptimal time allocation can lead to inefficient trajectories and underutilization of the quadrotor dynamic capabilities [3].
However, trajectory optimization methods often fail to fully leverage the quadrotor dynamic potential and struggle to operate effectively in high-agility scenarios [4]. Furthermore, the trajectories generated by these optimizers must be tracked by a controller, introducing delays, increased computational overhead, and mismatches between the optimized and actual flight trajectories, particularly under aggressive maneuvers.
@@
To address these challenges, we propose an improved Search-to-Control RL-based motion planner for quadrotor autonomous flight in dense environments. Our approach aims to maximize the utilization of the quadrotor dynamic capabilities while relying solely on a rate controller. Our method begins with a path searching algorithm based on heuristic search, which constructs a collision-free, minimum-length path in a discrete grid space (Fig. 2). This high-level path is then converted into control points, which serve as input for the RL control generator. The RL policy network processes observations of the quadrotor state, control points, and environmental information to generate low-level control commands. Finally, the generated commands are sent directly to the quadrotor rate loop controller, enabling the execution of agile and precise motions.
@@
In this letter, we design a Reinforcement LEarning based Search-to-Control framework called RESC. Compared to existing motion planners, our method departs from trajectory generation and optimization frameworks by directly generating control commands using a RL policy. This approach can handle discrete paths that may not inherently satisfy kinodynamic and non-holonomic constraints, producing aggressive, collision-free motions within specified dynamic limits, without relying on iterative numerical optimization. We evaluate the effectiveness and robustness of the proposed method through extensive simulations and real-world experiments across diverse scenarios. Our key contributions are summarized as follows:
A unified Search-to-Control planning framework that combines visibility-based path searching with RL-driven control generation, enabling aggressive and feasible motions without explicit trajectory representation or optimization.
A comprehensive RL environment that fully utilizes quadrotor dynamic potential to train policies, facilitating the generation of low-level control commands for diverse environments and seamless integration with the pathfinding module.
A reduced-dimensional obstacle observation method that efficiently captures environmental complexity, enabling lightweight neural network training and inference for robust control generation.
Extensive simulation and real-world experiments on an autonomous quadrotor, demonstrating the robustness and practical applicability of the proposed framework. The source code is released for further research.
@@ II Related Work
II-A Trajectory Optimization Methods
Trajectory optimization methods prevail as the predominant approach for quadrotor motion planning, which focus on generating smooth, dynamically feasible, and collision-free trajectories by modeling as optimal control problems. The pioneering approach is the minimum-snap trajectory generation[5] in differential flatness outputs, representing the trajectory as a polynomial and using gradient descent to iteratively modify the time segments. Several works propose two-stage solving methods where trajectory generation occurs in the second stage [6, 7, 8, 9, 10, 11]. For instance, a safe flight corridor composed of convex polyhedra is introduced in [6], providing constraints for optimization. Additionally, B-splines are utilized for kinodynamic path searching in [8], where an elastic optimization method is employed to refine trajectories. The Euclidean Signed Distance Field (ESDF) map is employed in [12] to obtain collision potentials and in [9] to find a path in the velocity field, improving time allocation. [10] proposes a comprehensive pipeline for quadrotor motion planning, including kinodynamic path searching, B-spline trajectory generation, time allocation, and nonuniform B-spline optimization based on ESDF. Subsequent works [13] and [14] introduce topological path and risk-aware yaw planning. Given the computational intensity of constructing ESDF, [15] proposes an ESDF-free gradient-based planner. Wang et al. [16] focuses on spatial-temporal deformation, which aligns with our approach by considering both geometric and temporal planning concurrently. These methods effectively generate smooth, safe, and feasible trajectories. However, they may be conservative regarding the ignorance of quadrotor dynamic capabilities.
@@ II-B RL based Methods
RL-based methods have gained significant traction in quadrotor motion planning due to their capability to optimize non-convex and discontinuous objectives. In [17], the authors propose an imitation learning (IL)-based teacher-student pipeline that directly maps noisy sensory depth images to polynomial trajectories. While IL is quite efficient, its scalability is limited by the need for expert data collection. The work in [18] utilizes a topological path-searching algorithm to generate guiding paths, which are subsequently refined by reinforcement learning to produce control commands. This two-stage method performs well in selected scenarios, outperforming some traditional methods. Song et al. [19] extracts features from depth images and uses a teacher-student IL framework similar to [17]. Furthermore, it generates control commands directly and includes a perception reward to guide yaw planning. In [20, 4], Song et al. investigates the strengths and limitations of optimal control and RL, proposing that RL is advantageous as it can optimize better objectives that may be infeasible for optimal control methods to address. In [21], Zhao et al. propose a hierarchical RL framework where an outer-loop policy dynamically configures speed and acceleration constraints based on online observations, significantly improving flight efficiency and safety in cluttered environments.
However, these methods frequently rely on imitation learning (IL) with expert knowledge or are trained in fixed environments, limiting their generalization and transferability to unknown environments, thereby hindering autonomous flight in new scenarios. Instead of using RL for the entire planning process, we employ it as an optimizer to refine control points and directly generate control commands. This strategy allows for more efficient learning and better adaptation to complex environments.
@@ III System Overview
III-A Planning Structure
The proposed framework, illustrated in Fig. 3, follows a planning process that consists of two stages. First, the Visibility Path Searching algorithm constructs a reference path from the start to the goal, ensuring obstacle avoidance and path length minimization. This reference path is discretized into control points based on the current position and constraints among control points (Section IV). Second, the RL Control Generation method is employed to generate control commands that optimize the quadrotor motion. These commands guide the quadrotor to follow the control points within a specified tolerance, determining the proximity requirements and strategies for obstacle avoidance in continuous space, as detailed in Section V.
III-B Quadrotor dynamics
The quadrotor is modelled as a rigid body governed by a linear system with decoupled translational and rotational dynamics. The control input vector corresponds to the individual rotor thrusts, specifically defined as with denotes the thrust scalar of the -th rotor. The complete state of the quadrotor is defined as , where is the position in world frame, represents roll, pitch and yaw angle, and denotes angular velocity in body frame.
The dynamics equations are
where is velocity, is rotation matrix from body frame to world frame, is z axis in body frame, is thrust of each rotor, is mass, is diagonal inertia matrix, , and is gyroscopic torque and torque by thrusts respectively. We use Euler angles to represent rotation and set a limit on roll and pitch angles to avoid singularity.
@@ III-B Quadrotor dynamics
where denotes rotor’s torque coefficient, and represents the distance from the center of mass to the rotor center.
@@ IV Visibility Path Searching
Our path searching algorithm combines elements of the A*, jump point search (JPS) [22], and visibility graph methods. We inherit the heuristic function and the open and closed list architecture from the A* algorithm but implement an alternative cost function. Elements of the JPS method are incorporated to reduce the number of nodes in the open list, minimizing computational costs associated with sorting these nodes. Additionally, we leverage the optimal distance characteristics of the visibility graph method without constructing complex polygon edges and vertices.
The proposed strategy facilitates omnidirectional movement through obstacle corner detection and visibility checks, rather than restricting movement to increments of in 2D plane. This approach is implemented by selecting obstacle corners in the Corner function and assessing voxel visibility in the VisibleCheck function. As a result, this method reduces the computational and memory overhead typically required for accessing each vertex and edge of the voxels. Subsequently, the corners are linked to form paths, which are then sampled to generate control points. The detailed algorithm is presented in Algorithm 1.
@@ V-A Policy Framework
We frame the control point optimization and control generation problem within a reinforcement learning (RL) framework. The quadrotor agent selects actions based on the policy , transitioning from state to according to the dynamics-kinematics update of the system. The agent receives a reward that guides learning. Detailed configurations of the state space, action space, and reward function will be discussed in subsequent sections.
@@ V-A1 Observation and action space
The observation space contains three main components: the state of the quadrotor, control points and obstacle information. The state vector is defined as
where represents control points, and denotes obstacle information, .
The observation vector is derived from the state vector, directly incorporating components such as , , , and . Other elements of the state are converted into lower-level features, making them more generalized and easier for policy training. For example, we use relative position from quadrotor position to control points instead of absolute position. Additionally, a normalized vector representing the x-axis direction of body frame in world frame () is used to guide the yaw to stay visiualized for control point based on the field-of-view (FOV) of depth camera. Obstacle observation will be discussed in the following section.
The action produced by RL policy is , where is the collective thrust and represents the desired body rate. This control approach has been proved effective for learning-based quadrotor control and planning tasks [25]. Although commonly used in drone racing tasks due to its low latency and high frequency, it can present a gap between simulation and real-world performance. Despite this limitation, this control approach leads to smoother acceleration and velocity, improving energy efficiency and safety compared to higher-dimensional control approaches.
@@ V-A2 Reward function
According to paper[20], RL outperforms Optimal Control (OC) in terms of optimizing a better objective that OC may not be able to solve. From this perspective, we design a reward function that explicitly represents agile motion, collision avoidance, and smoothness in control. The reward function is defined as
where denotes the customized progress reward, represents the penalty for collisions, is the penalty for dynamic violations, is the penalty for FOV violations, and rewards task completion.
@@
where denotes the intersection point if the vector intersects the control point area; otherwise, the intersection point is directly . We normalize the progress reward and subsequently introduce a negative term , thereby imposing a time penalty which maximizing the utilization of the quadrotor dynamic capabilities.
For dynamic violation penalty , we impose limits on acceleration and velocity to ensure dynamic smoothness and safety. Additionally, attitude limits are set to prevent overly aggressive maneuvers and avoid singularities.
where represents the number of elements from the set that exceed their prescribed limits.
The FOV violation penalty is set to if neither of the next two control points is within the FOV of the depth camera. This ensures that the quadrotor maintains visibility of the control points.
@@ V-B Training methodology
We create a training environment based on OpenAI gym[26], providing an interface for creating a customized environment. At the beginning of each episode, the velocity, body rate, orientation, control points and obstacles are randomly initialized following a predefined probabilistic pattern, while the position if fixed at the center of the map (Fig. 4).
V-B1 Environment randomization
The environment is initialized with control points and obstacles, with the quadrotor starting at the map center. The first control point is randomized within a specified distance from the quadrotor position. Subsequent control points are placed within a defined distance and angle from the previous one. Velocity and orientation constraints are applied to keep the quadrotor motion stable, ensuring its velocity remains basically towards the first control point, and its pitch, roll, and yaw stay within specified limits.
Obstacles are placed along or near the lines connecting the quadrotor to the control points, ensuring collision-free paths while allowing some proximity to obstacles, rather than creating a maze-like environment.
V-B2 Control point observation
Following Song et al. [4], we observe two relative positions of the control points. For the last control point, the relative position of the quadrotor to this point is repeated. We define a finite-height column with a radius of and a height of to represent the area within which the control point is considered passed. The observation is then updated to the next two control points. With this design and the randomization of the environment, the use of a small number of control points (we use two in our training) enables efficient training while allowing generalization to more complex tasks with a larger number of control points.
V-B3 Obstacle observation
We introduce a novel method to describe obstacle information efficiently, called Kinematic Guided Pseudo-Raycast and ESDF Perception (KGP-ESDF), as shown in Fig. 5.
The core concept of KGP is to identify the most relevant obstacle edges for the quadrotor motion towards the control point, simultaneously avoiding collisions along the current velocity direction. First, the algorithm performs a sector-based search along the line from the current position to the control point to identify the nearest obstacle edges, denoted as . Subsequently, a similar search is conducted along the line from to the velocity control point as defined by the equation
to identify the nearest obstacle edges, denoted as .
The relative positions of these edge points are incorporated as part of the observation. Additionally, we record the distance from the current position to the velocity control point, which represents the collision-free distance.
We generate to represent the surrounding obstacle information, derived from the ESDF map. Finally, the obstacle observation is constructed as
V-C Sim-to-real transfer